OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (mertsj):

What is the slope of the given line?

OpenStudy (anonymous):

i'm not really sure but these are my answer choices: y = –x + 5 y = –x – 1 y = x + 1 y = x + 5

OpenStudy (mertsj):